# Measuring What Matters: The AI Pluralism Index

### v1

@@ 1 Introduction
Artificial intelligence is now infrastructural in search, education, media, health, finance, and public administration. Decisions about data curation, model objectives, deployment defaults, safety mitigations, and redress are largely made by a small number of labs and government programs. International guidance warns that such concentration undermines democratic legitimacy and risks sustained exclusion of those most affected (UNESCO, 2021; Crawford, 2021; Benkler, 2006; Council of Europe, 2024). Users face a second challenge: for many everyday tasks, top systems often produce broadly similar outputs; headline capability differences are therefore less informative for responsible selection (Ray, 2023; Bommasani et al., 2022; Chiang et al., 2024; Stanford Center for Research on Foundation Models (2025), CRFM). When functionality converges, the governance properties of systems become a salient differentiator.
Existing evaluations compare model capabilities and user preferences (Ray, 2023; Chiang et al., 2024; Stanford CRFM, 2024). Other efforts synthesize ethics and disclosure guidance (Jobin et al., 2019; Hagendorff, 2020; Mushkani et al., 2025b; Bommasani et al., 2023; Iyer, 2024) and safety governance (Bengio et al., 2024; Future of Life Institute, 2025; Tabassi et al., 2023; European Union, 2024a; International Organization for Standardization, 2023). Regulatory frameworks articulate process expectations—risk management, incident handling, documentation, accessibility—but do not themselves yield comparable scores across providers (Tabassi, 2023; European Union, 2024b). There is no comprehensive, public, auditable instrument that centers pluralism—whether affected stakeholders can influence objectives, safeguards, and deployment—and measures participation, inclusion, transparency, and accountability in verifiable ways. This paper addresses that gap.
@@
This paper introduces the AI Pluralism Index (AIPI), a community‑governed protocol that: (i) formalizes an indicator hierarchy operationalizing pluralism across four pillars grounded in international norms and empirical work (Jobin et al., 2019; Arnstein, 1969; Bohman, 2000; Mitchell et al., 2019; Gebru et al., 2021; Goggin et al., 2019; UNESCO, 2021; OECD, 2019; Fung, 2006; World Wide Web Consortium (2023), W3C); (ii) implements a transparent evidence pipeline and scoring with two principled treatments of Unknown; (iii) evaluates reliability and construct validity via inter‑rater agreement, cross‑index correlations, and sensitivity analysis; and (iv) publishes system‑ and provider‑level results with evidence provenance through a versioned, open governance process aligned with standards for incident reporting and vulnerability disclosure (International Organization for Standardization, 2018, 2019; Responsible AI Collaborative, 2024).
@@ 2 Related work
Capability benchmarks such as Chatbot Arena and HELM provide large‑scale comparative results for tasks under controlled protocols (Ray, 2023; Chiang et al., 2024; Stanford CRFM, 2024). These efforts inform model selection for performance but are largely orthogonal to governance. Transparency syntheses assess disclosure practices and ethics guidance (Jobin et al., 2019; Hagendorff, 2020; Bommasani et al., 2023; Iyer, 2024). Safety indices and governance assessments focus on risk management commitments and safeguards (Bengio et al., 2024; Future of Life Institute, 2025). Regulatory instruments set expectations for trustworthy AI without creating measurement tools: the EU AI Act defines risk‑based obligations, documentation, and oversight (European Union, 2024b). The NIST AI Risk Management Framework provides organizational process scaffolding (Tabassi, 2023; Tabassi et al., 2023; International Organization for Standardization, 2023). AIPI complements these efforts by offering an auditable, evidence‑driven index focused on pluralism: participation, inclusion, transparency, and accountability practices that shape how systems are built and governed.
@@ 3 Concept and scope
We define AI pluralism as a property of socio‑technical development and deployment processes whereby affected stakeholders can meaningfully shape objectives, constraints, and outcomes. Pluralism concerns both governance and production. The index is anchored in four equal‑weighted pillars (Fig. 1): participatory governance, inclusivity and diversity, transparency, and accountability. These pillars align with traditions in participation studies and human–computer interaction—Arnstein’s ladder and deliberative traditions emphasize degrees of influence and institutional design (Arnstein, 1969; Bohman, 2000; Fung, 2006); Value Sensitive Design operationalizes stakeholder engagement and value articulation in technology development (Umbrello and van de Poel, 2021; Friedman and Hendry, 2019). Inclusivity requires attention to representation in teams, language access, and accessibility (Goggin et al., 2019; World Wide Web Consortium (2023), W3C). Transparency and accountability cover documentation, incident reporting, audits, redress, and vulnerability handling (Mitchell et al., 2019; Gebru et al., 2021; Lepri et al., 2018; International Organization for Standardization, 2018, 2019). The unit of analysis is the producer (company, lab, consortium, or public agency), with optional system‑family granularity when practices differ across offerings. The scope is global and normative, aligning with recognized principles rather than a single jurisdiction.
@@ 4 Design choices: indicators, observability, and consequences
AIPI credits only practices that leave public, auditable traces. Indicators must be observable through public artifacts or external evaluations, verifiable via durable references and timestamps (with archives when permitted), and consequential in the sense that they allocate decision rights, impose process requirements, or document changes. This approach reduces coder discretion, enables third‑party audit, and aligns scoring with practices that affected stakeholders can inspect and use.
@@
Verifiability requires that evidence be durable and comparable across releases. Each artifact must have a persistent URL and a clear date; where possible we store or reference an archived copy. Versioned artifacts are coded by version. Materials behind paywalls, NDAs, or logins do not qualify as primary evidence because affected stakeholders cannot inspect them. When a standards claim is made (e.g., “WCAG 2.2 AA”), the artifact should state scope (surfaces, products, versions) and date; claims without scope or date receive partial credit under the ordinal rubric in Table 1 (Goggin et al., 2019; World Wide Web Consortium (2023), W3C).
@@
For inclusivity and diversity, we count language support only when multiple user‑facing components are present for the flagship surface. A language is counted if at least three of the following are documented: localized user interface, localized help or FAQ, localized policies and safety disclosures, and a localized redress or appeals channel. Auto‑translation or preview‑only modes are not counted. Accessibility claims are assessed with stated scope and date; self‑declarations without scope receive partial credit; independent verification receives full credit (Goggin et al., 2019; World Wide Web Consortium (2023), W3C). Partnerships outside high‑income markets count when they involve co‑production (e.g., co‑authored resources, shared maintenance) rather than one‑off donations. Statements of commitment without accompanying statistics, methods, or changes do not receive credit. For multilingual evaluation and collaboration in lower‑resourced settings, we draw on evidence from participatory language work (Nekoto et al., 2020).
@@
Where mature standards exist, coding is anchored to them to reduce ambiguity across providers and releases. For accessibility we assess the presence of concrete conformance claims and independent verification (Goggin et al., 2019); for security disclosure and handling we look for documented processes (Lepri et al., 2018; Leslie, 2019). Standards references without scope or date do not suffice; conformance statements with scope and date receive partial credit; verification (e.g., audit or test report) receives full credit (World Wide Web Consortium (2023), W3C; International Organization for Standardization, 2018, 2019).
@@ 5 Measurement model
AIPI aggregates evidence‑coded indicators to pillar and overall scores with explicit treatment of Unknown. Figure 2 summarizes the pipeline; Figure 3 shows the pillar–subpillar–indicator hierarchy.
@@ 6 Indicators and coding guide
Indicators are designed to be observable, verifiable, and consequential. Each code must be supported by a link to a public artifact (policy, governance record, model/system card, audit report, standard conformance statement) or a vetted external evaluation. Detailed rubrics and edge cases are in the public codebook (anonymized for peer review). Table 1 summarizes the generic ordinal mapping used when indicators admit partial credit.
@@
This pillar codes whether affected stakeholders can provide input that is documented, deliberated, and linked to resulting actions. Qualifying evidence consists of (i) multi‑stakeholder advisory bodies with a public charter defining remit, membership selection, conflict‑of‑interest handling, and publication of agendas and minutes; (ii) public consultations or request‑for‑comment processes that publish submissions in aggregate and a disposition matrix mapping proposals to accept/reject/rationale; (iii) issue trackers or proposal forums open to external contributors where each proposal has a dated, public disposition and, when adopted, a pointer to the corresponding change log or release note; (iv) external community red teaming with a published scope, methodology, consent and safety procedures, and a post‑exercise report that identifies specific changes or risk acceptances; and (v) commitments to avoid undue concentration of power that are evidenced by implemented access or governance rules (e.g., non‑exclusive licensing terms, neutral API access criteria, data export or interoperability guarantees, limits on exclusive partnerships) rather than aspirational statements (Arnstein, 1969; Bohman, 2000; Open Government Partnership, 2024; Fung, 2006). Credit is assigned to artifacts that are publicly accessible at or before the release cutoff and that show at least one completed feedback cycle (input, deliberation, decision, implementation or reasoned rejection). Ad hoc surveys without published dispositions, private advisory arrangements without a public mandate, and announcements without links to decisions do not qualify (Mushkani et al., 2025a).
@@ 8 Data sources and pipeline
The pipeline integrates three streams. First, structured web and repository search traverses producer sites, documentation hubs, transparency reports, model and dataset cards, and registries; enumeration uses Google search and curated news articles to seed candidate providers, then applies de‑duplication and eligibility filters (Sec. 7). Reproducibility metadata include query templates, URLs, timestamps, and retrieval logs. Second, external independent evaluations and benchmarks are incorporated with explicit citation, for example large‑scale capability benchmark syntheses and public safety assessments (Ray, 2023; Bengio et al., 2024). Third, expert interviews with researchers, journalists, and civil society practitioners provide context for interpreting artifacts and identifying missing documentation; interviews obtain informed consent and exclude sensitive data. Evidence links are checked for accessibility and archived when permitted. The evidence graph ties producers to indicators, sources, and coders; adjudication logs are public (anonymized for review). All builds are deterministic and versioned.
@@ 11 Governance, implementation, and releases
AIPI is implemented as a public repository with deterministic builds and archived artifacts (anonymized for review). Anyone can file an issue to challenge a code, propose an indicator change, or contribute evidence. Issues are triaged by volunteer reviewers; disagreements are logged and escalated to a multi‑stakeholder maintainer council spanning research, civil society, industry, and regions. Each release is versioned and archived with a change log. The website provides system and provider profiles, an API, and download links for evidence tables and scores. Governance adheres to the same pluralistic principles that the instrument evaluates and aligns with open practices in accessibility and security communities (Goggin et al., 2019; Lepri et al., 2018).
@@ 12 Recommendations for practice and policy
For producers, a practical path to improving AIPI is to publish verifiable artifacts that reflect real processes. Model and system cards should disclose data governance, known limitations, intended uses, subgroup performance, and update histories. Vulnerability disclosure processes should accept reports from good‑faith researchers and specify service levels; vulnerability handling should name a PSIRT function and lay out triage and remediation steps. Public consultations should summarize input and document how it influenced decisions rather than collecting comments without response. Disaggregation in evaluation should be routine for quality and safety. For procurers, reference the AIPI indicator list directly in requests for proposals, require minimum evidence scores and explicit coverage thresholds, and use pillar‑level scores to tailor due‑diligence questions. For regulators, several indicators map to policy requirements under the EU AI Act and to current governance debates—risk management, incident reporting, documentation, and redress—and can be incorporated as audit checks (European Union, 2024b; Ulnicane, 2024). For researchers and civil society, AIPI’s evidence graph enables systematic scrutiny of claims and targeted campaigns that reward concrete improvements.

### v3

@@ 1 Introduction
Artificial intelligence is now infrastructural in search, education, media, health, finance, and public administration. Decisions about data curation, model objectives, deployment defaults, safety mitigations, and redress are largely made by a small number of labs and government programs. International guidance warns that such concentration undermines democratic legitimacy and risks sustained exclusion of those most affected (UNESCO, 2021; Crawford, 2021; Benkler, 2006; Council of Europe, 2024). Users face a second challenge: for many everyday tasks, top systems often produce broadly similar outputs; headline capability differences are therefore less informative for responsible selection (Ray, 2023; Bommasani et al., 2022; Chiang et al., 2024; Stanford Center for Research on Foundation Models (CRFM), 2025). When functionality converges, the governance properties of systems become a salient differentiator.
Existing evaluations compare model capabilities and user preferences (Ray, 2023; Chiang et al., 2024; Stanford CRFM, 2024). Other efforts synthesize ethics and disclosure guidance (Jobin et al., 2019; Hagendorff, 2020; Mushkani et al., 2025; Bommasani et al., 2023; Iyer, 2024) and safety governance (Bengio et al., 2024; Future of Life Institute, 2025; European Union, 2024a; International Organization for Standardization, 2023). Regulatory frameworks articulate process expectations—risk management, incident handling, documentation, accessibility—but do not themselves yield comparable scores across providers (Tabassi, 2023; European Union, 2024b). There is no comprehensive, public, auditable instrument that centers pluralism—whether affected stakeholders can influence objectives, safeguards, and deployment—and measures participation, inclusion, transparency, and accountability in verifiable ways. This paper addresses that gap.
@@
This paper introduces the AI Pluralism Index (AIPI), a community‑governed protocol that: (i) formalizes an indicator hierarchy operationalizing pluralism across four pillars grounded in international norms and empirical work (Jobin et al., 2019; Arnstein, 1969; Bohman, 2000; Mitchell et al., 2019; Gebru et al., 2021; Goggin et al., 2019; UNESCO, 2021; OECD, 2019; Fung, 2006; World Wide Web Consortium (W3C), 2023); (ii) implements a transparent evidence pipeline and scoring with two principled treatments of Unknown; (iii) evaluates reliability and construct validity via inter‑rater agreement, cross‑index correlations, and sensitivity analysis; and (iv) publishes system‑ and provider‑level results with evidence provenance through a versioned, open governance process aligned with standards for incident reporting and vulnerability disclosure (International Organization for Standardization, 2018, 2019; Responsible AI Collaborative, 2024).
@@ 2 Related work
Capability benchmarks such as Chatbot Arena and HELM provide large‑scale comparative results for tasks under controlled protocols (Ray, 2023; Chiang et al., 2024; Stanford CRFM, 2024). These efforts inform model selection for performance but are largely orthogonal to governance. Transparency syntheses assess disclosure practices and ethics guidance (Jobin et al., 2019; Hagendorff, 2020; Bommasani et al., 2023; Iyer, 2024). Safety indices and governance assessments focus on risk management commitments and safeguards (Bengio et al., 2024; Future of Life Institute, 2025). Regulatory instruments set expectations for trustworthy AI without creating measurement tools: the EU AI Act defines risk‑based obligations, documentation, and oversight (European Union, 2024b). The NIST AI Risk Management Framework provides organizational process scaffolding (Tabassi, 2023; International Organization for Standardization, 2023). AIPI complements these efforts by offering an auditable, evidence‑driven index focused on pluralism: participation, inclusion, transparency, and accountability practices that shape how systems are built and governed.
@@ 3 Concept and scope
We define AI pluralism as a property of socio‑technical development and deployment processes whereby affected stakeholders can meaningfully shape objectives, constraints, and outcomes. Pluralism concerns both governance and production. The index is anchored in four equal‑weighted pillars (Fig. 1): participatory governance, inclusivity and diversity, transparency, and accountability. These pillars align with traditions in participation studies and human–computer interaction—Arnstein’s ladder and deliberative traditions emphasize degrees of influence and institutional design (Arnstein, 1969; Bohman, 2000; Fung, 2006); Value Sensitive Design operationalizes stakeholder engagement and value articulation in technology development (Umbrello and van de Poel, 2021; Friedman and Hendry, 2019). Inclusivity requires attention to representation in teams, language access, and accessibility (Goggin et al., 2019; World Wide Web Consortium (W3C), 2023). Transparency and accountability cover documentation, incident reporting, audits, redress, and vulnerability handling (Mitchell et al., 2019; Gebru et al., 2021; Lepri et al., 2018; International Organization for Standardization, 2018, 2019). The unit of analysis is the producer (company, lab, consortium, or public agency), with optional system‑family granularity when practices differ across offerings. The scope is global and normative, aligning with recognized principles rather than a single jurisdiction. Because AIPI is intended as a global index, participatory governance indicators distinguish between geographically bounded engagement (e.g., only in a producer’s home jurisdiction) and engagement that is demonstrably multi-region. We do not apply separate region or population weights in aggregation; instead, geographic scope is captured within relevant indicator rubrics and reported as metadata in the evidence graph.
@@ 4 Design choices: indicators, observability, and consequences
AIPI credits only practices that leave public, auditable traces. Indicators must be observable through public artifacts or external evaluations, verifiable via durable references and timestamps (with archives when permitted), and consequential in the sense that they allocate decision rights, impose process requirements, or document changes. This approach reduces coder discretion, enables third‑party audit, and aligns scoring with practices that affected stakeholders can inspect and use. More broadly, AIPI follows established methodology for constructing composite indicators, including explicit conceptual framing, transparent normalization and weighting choices, and sensitivity analysis (OECD et al., 2008).
@@
Verifiability requires that evidence be durable and comparable across releases. Each artifact must have a persistent URL and a clear date; where possible we store or reference an archived copy. Versioned artifacts are coded by version. Materials behind paywalls, NDAs, or logins do not qualify as primary evidence because affected stakeholders cannot inspect them. When a standards claim is made (e.g., “WCAG 2.2 AA”), the artifact should state scope (surfaces, products, versions) and date; claims without scope or date receive partial credit under the ordinal rubric in Table 1 (Goggin et al., 2019; World Wide Web Consortium (W3C), 2023).
@@
For inclusivity and diversity, we count language support only when multiple user‑facing components are present for the flagship surface. A language is counted if at least three of the following are documented: localized user interface, localized help or FAQ, localized policies and safety disclosures, and a localized redress or appeals channel. Auto‑translation or preview‑only modes are not counted. Accessibility claims are assessed with stated scope and date; self‑declarations without scope receive partial credit; independent verification receives full credit (Goggin et al., 2019; World Wide Web Consortium (W3C), 2023). Partnerships outside high‑income markets count when they involve co‑production (e.g., co‑authored resources, shared maintenance) rather than one‑off donations. Statements of commitment without accompanying statistics, methods, or changes do not receive credit. For multilingual evaluation and collaboration in lower‑resourced settings, we draw on evidence from participatory language work (Nekoto et al., 2020).
@@
Where mature standards exist, coding is anchored to them to reduce ambiguity across providers and releases. For accessibility we assess the presence of concrete conformance claims and independent verification (Goggin et al., 2019); for security disclosure and handling we look for documented processes (Lepri et al., 2018; Leslie, 2019). Standards references without scope or date do not suffice; conformance statements with scope and date receive partial credit; verification (e.g., audit or test report) receives full credit (World Wide Web Consortium (W3C), 2023; International Organization for Standardization, 2018, 2019).
@@ 5 Measurement model
AIPI aggregates evidence-coded indicators into pillar and overall scores with explicit treatment of Unknown. Figure 2 provides a visual overview of the scoring logic presented in this section. Figure 3 summarizes the overall pipeline, and Figure 4 presents the pillar–subpillar–indicator hierarchy.
Figure 2: Simple schematic of the AIPI measurement model. Indicators are coded from verifiable artifacts, normalized, aggregated to pillar means, and reported as an interval via explicit treatments of Unknown (evidence, known-only, optimistic) with coverage.
@@ 6 Indicators and coding guide
Indicators are designed to be observable, verifiable, and consequential. Each code must be supported by a link to a public artifact (policy, governance record, model/system card, audit report, standard conformance statement) or a vetted external evaluation. Detailed rubrics and edge cases are available in the public codebook on GitHub:
Table 1 summarizes the generic ordinal mapping used when indicators admit partial credit.
@@
This pillar codes whether affected stakeholders can provide input that is documented, deliberated, and linked to resulting actions. Qualifying evidence consists of (i) multi‑stakeholder advisory bodies with a public charter defining remit, membership selection, conflict‑of‑interest handling, and publication of agendas and minutes; (ii) public consultations or request‑for‑comment processes that publish submissions in aggregate and a disposition matrix mapping proposals to accept/reject/rationale; (iii) issue trackers or proposal forums open to external contributors where each proposal has a dated, public disposition and, when adopted, a pointer to the corresponding change log or release note; (iv) external community red teaming with a published scope, methodology, consent and safety procedures, and a post‑exercise report that identifies specific changes or risk acceptances; and (v) commitments to avoid undue concentration of power that are evidenced by implemented access or governance rules (e.g., non‑exclusive licensing terms, neutral API access criteria, data export or interoperability guarantees, limits on exclusive partnerships) rather than aspirational statements (Arnstein, 1969; Bohman, 2000; Open Government Partnership, 2024; Fung, 2006). Because AIPI is a global index, we also code the geographic scope of participatory mechanisms when evidence permits. For advisory bodies, public consultations, and community red teaming, full credit (1.0) requires evidence of participation that is multi-region (e.g., a roster or contributor list spanning multiple world regions, or an open process accessible internationally, including across languages where relevant). Partial credit (0.5) is assigned when a mechanism exists but participation appears limited to a single country/region, or when the geographic composition is not disclosed; no credit (0.0) is assigned when no verifiable mechanism exists.
Credit is assigned to artifacts that are publicly accessible at or before the release cutoff and that show at least one completed feedback cycle (input, deliberation, decision, implementation or reasoned rejection). Ad hoc surveys without published dispositions, private advisory arrangements without a public mandate, and announcements without links to decisions do not qualify.
@@ 8 Data sources and pipeline
The pipeline integrates three streams. First, structured web and repository search traverses producer sites, documentation hubs, transparency reports, model and dataset cards, and registries; enumeration uses Google search and curated news articles to seed candidate providers, then applies de‑duplication and eligibility filters (Sec. 7). Reproducibility metadata include query templates, URLs, timestamps, and retrieval logs. Second, external independent evaluations and benchmarks are incorporated with explicit citation, for example large‑scale capability benchmark syntheses and public safety assessments (Ray, 2023; Bengio et al., 2024). Third, expert interviews with researchers, journalists, and civil society practitioners provide context for interpreting artifacts and identifying missing documentation; interviews obtain informed consent and exclude sensitive data. Evidence links are checked for accessibility and archived when permitted. The evidence graph ties producers to indicators, sources, and coders; adjudication logs are public. All builds are deterministic and versioned.
@@ 11 Governance, implementation, and releases
AIPI is implemented as a public repository with deterministic builds and archived artifacts. Anyone can file an issue to challenge a code, propose an indicator change, or contribute evidence. Issues are triaged by volunteer reviewers; disagreements are logged and escalated to a multi‑stakeholder maintainer council spanning research, civil society, industry, and regions. Each release is versioned and archived with a change log. The website provides system and provider profiles, an API, and download links for evidence tables and scores. Governance adheres to the same pluralistic principles that the instrument evaluates and aligns with open practices in accessibility and security communities (Goggin et al., 2019; Lepri et al., 2018).
We operationalize this governance with three layers of decision making. First, a release team executes evidence discovery, coding, and quality assurance using a frozen codebook for that version. Second, an adjudication panel resolves disputed codes using a publicly documented standard of evidence, logging rationales and precedents so that similar cases are treated consistently across releases. Third, the maintainer council decides on indicator additions, deprecations, and major rubric changes through a transparent proposal process that includes (i) an open call for change requests, (ii) a public comment window, and (iii) recorded decisions with an explicit migration plan for longitudinal comparability.
To address concentration and geographic skew in participation, the maintainer council is constituted to include representation across stakeholder categories and regions (research, civil society, industry, labor, and policy), with time-bounded terms and conflict-of-interest disclosures. Members recuse themselves from decisions involving their own organizations. Releases follow a predictable cadence with an evidence freeze date, double-coding of a stratified sample, and a short pre-release audit period during which external contributors can contest evidence links or propose missing artifacts. Final releases are archived, and we publish a changelog that distinguishes schema changes (which affect comparability) from evidence updates (which affect scores).
@@ 12 Recommendations for practice and policy
For producers, a practical path to improving AIPI is to publish verifiable artifacts that reflect real processes, and to ensure that those processes include affected stakeholders across deployment contexts. Model and system cards should disclose data governance, known limitations, intended uses, subgroup performance, and update histories. Vulnerability disclosure processes should accept reports from good-faith researchers and specify service levels; vulnerability handling should name a PSIRT (Product Security Incident Response Team) function and lay out triage and remediation steps. Public consultations should summarize input and publish a disposition matrix that links proposals to decisions and downstream changes, rather than collecting comments without response. Where systems are deployed internationally, participatory mechanisms should be accessible across regions, for example via remote participation, multilingual materials, and geographically diverse advisory membership.
For procurers, AIPI can be used as a practical checklist rather than a headline ranking. Reference the indicator list directly in requests for proposals; require minimum evidence scores and explicit coverage thresholds; and ask vendors to supply the primary artifacts (not summaries) referenced by their scores. Use pillar-level scores to tailor due-diligence: for low participatory-governance scores, require a roadmap for consultation and redress; for low transparency scores, require current system cards and change logs; for low accountability scores, require incident response and audit documentation.
For regulators and standards bodies, several indicators map onto emerging governance requirements, including documentation, risk management, incident reporting, accessibility, and redress. AIPI can support supervision by translating these requirements into auditable evidence checks and by highlighting where documentation is absent, stale, or geographically narrow. Regulators can also use AIPI’s coverage reporting to prioritize oversight, focusing audits where low scores are paired with low evidence coverage.
For researchers and civil society, AIPI’s evidence graph enables systematic scrutiny of claims and targeted advocacy that rewards concrete improvements. We encourage third-party replications and “shadow scores” for contested providers, and we treat well-documented replications as admissible evidence in subsequent releases.